The Kincaid Company uses a predetermined overhead rate to apply manufacturing overhead to production. The rate is based on direct labor hours. Estimates for the year just ended are as follows:

Estimated Manufacturing Overhead $240,000

Estimated Direct Labor Hours 40,000

At the end of the year, Kincaid Company records revealed the following information:

Raw Materials Inventory $35,000

Work-In-Process Inventory 60,000

Finished Goods inventory 105,000

Cost of Goods Sold 400,000

Manufacturing Overhead costs incurred 210,000

Actual Direct Labor Hours 39,000

Required:

a. Calculate the predetermined overhead rate for the year.

b. Determine the amount of overhead applied during the year.

c. Determine the amount of underapplied/overapplied manufacturing overhead for the year.
Answer #1

a) Predetermine overhead rate = 240000/40000 = 6 per DLH

b) Applied overhead = 39000*6 = 234000

c) Actual manufacturing overhead = 210000

Applied manufacturing overhead = 234000

Over applied overhead = 234000-210000 = 24000